Suppose two bees, each with a charge of 93.7 pC, are separated by a distance of 1.21 cm

1.Suppose two bees, each with a charge of 93.7 pC, are separated by a distance of 1.21 cm. Treating the bees as point charges, with the magnitude of the electrostatic force experienced by the bees. Compare this force to the weight of a 0.112 g bee. By what factor is the force smaller than the bee? 
Answer #1

given charge of each bee=93.7 pc

1pc=10-12coloumbs

distance between two bees=1.21cm=1.21/100=0.0121m

According to Columb's law F=1/4πε0 * Q1*Q2/r2=[(9.0 x 109 N•m2/C2)*(93.7*10-12coloumbs)*(93.7*10-12coloumbs)] / (0.0121m)2 F=0.0000005397=5397*10-10Newtons

here Q1 and Q2 are charges of bee, and r is the distance between the bee's

weight of bee with mass 0.122g=mass*g=0.122*9.8m/s2*10-3=1.1956*10-3Newtons

weight of the bee is greater than the electrostatic force between the two bees because the gravitational pull is greater than the force of attraction between two charged bees.


answered by: NEC